An element with 27gm mol^-1 forms a cubic unit cell with edge length 4.05*(10^-8) cm. If its density is 2.7gm cm^-3. what is the nature of the cubic unit cell l?

Grade:12

1 Answers

Arun
5 years ago

It is given that density of the element, d = 2.7 ×103kg m-3

Molar mass, M = 2.7 ×10-2 kg mol-1

Edge length, a= 405 pm = 405 ×10-12 m = 4.05 ×10-10 m

It is known that, Avogadro's number, NA= 6.022 ×1023 mol-1

d = zm /Na * a³

Applying the relation, This implies that four atoms of the element are present per unit cell. Hence, the unit cell is face-centred cubic (fcc) or cubic close-packed (ccp).